Solve the formula for the given variable. Assume all variables represent nonnegative numbers.
A=5s^2, for s.

Answer by edjones(7569)

5s^2=A
s^2=A/5
s=+-sqrt(A)/sqrt(5) Sqrt of both sides.
s=+-sqrt(A)/sqrt(5) * sqrt(5)/sqrt(5)
s=+-sqrt(5A)/5
s=sqrt(5A)/5, s=-sqrt(5A)/5
.
Ed